OpenStudy (anonymous):

how to solve for y? \[(x-y)(x-y)=9\]

OpenStudy (unklerhaukus):

\[(x-y)(x-y)=9\] \[(x-y)^2=3^2\] \[\qquad=\]

OpenStudy (tyteen4a03):

You cannot solve for y because there are 2 variables yet there are only 1 equation. Do you mean change of subject?

OpenStudy (lgbasallote):

here comes the race...

OpenStudy (tyteen4a03):

@Igbsallote those are fun

OpenStudy (anonymous):

so then sqrt and get x-y=3 and so y = x-3

OpenStudy (anonymous):

and y=x+3 also

OpenStudy (anonymous):

yes? :D

OpenStudy (unklerhaukus):

\[\LARGE\color{red}{\checkmark\checkmark}\]

OpenStudy (anonymous):

okay can you help me factor \[x^2-3xy-4y^2=0\]

OpenStudy (anonymous):

(x+y)(x-4y)?

hartnn (hartnn):

\(\LARGE\color{red}{\checkmark\checkmark}\)

OpenStudy (anonymous):

lol okay so if i was to solve that for y i would get... x+y=4y+x now what? :|

hartnn (hartnn):

(x+y)(x-4y) =0 x+y= 0 or x-4y = 0

OpenStudy (anonymous):

ooohhh :O okay so y=x or y=-1/4x

OpenStudy (anonymous):

y=-x*

hartnn (hartnn):

x+y=0 y=-x x-4y=0 y= x/4

OpenStudy (anonymous):

yup yup...col i was right ^_^

OpenStudy (anonymous):

cool* ...wow lol

hartnn (hartnn):

next question...

OpenStudy (lgbasallote):

someone took a detour and suddenly appeared in the race out of nowhere

OpenStudy (anonymous):

lol...ill post the next one when i have one..thanks guys! :)

hartnn (hartnn):

its no race for me

hartnn (hartnn):

welcome ^_^

OpenStudy (lgbasallote):

that's what they all say

OpenStudy (anonymous):

lol wut?

OpenStudy (tyteen4a03):

@lgbasallote Oh the irony :P

OpenStudy (lgbasallote):

where's the irony?

OpenStudy (tyteen4a03):

@lgbasallote Do I need to be that obvious? *wink*

OpenStudy (lgbasallote):

im not racing with any of you....

OpenStudy (lgbasallote):

i don't reply when someone else is replying

OpenStudy (anonymous):

@tyteen4a03 ur tryin way to hard buddy

OpenStudy (anonymous):

lol okay one last one...you guys have 4 minutes to explain it to me

OpenStudy (anonymous):

\[{3x+5}\over{6y-1}\] solve for y

OpenStudy (anonymous):

3y*

OpenStudy (unklerhaukus):

thats not an equation

OpenStudy (lgbasallote):

will the equation self-destruct if they cant solve it in 4 minutes?

OpenStudy (anonymous):

okay lemme write that again... \[x={{3y-5}\over{6y-1}}\]

OpenStudy (anonymous):

solve for y.....and i leave in 3 minutes thats why

OpenStudy (anonymous):

its +5...sorry

OpenStudy (anonymous):

3y+5

OpenStudy (unklerhaukus):

multiply by the denominator

OpenStudy (tyteen4a03):

@yummydum I'm trying what? :P

OpenStudy (anonymous):

that gives \[{18y^2+29y-5}\over {36y^2+12y+1}\]

OpenStudy (unklerhaukus):

agh, not what i ment

OpenStudy (anonymous):

:\

OpenStudy (anonymous):

im confused

OpenStudy (unklerhaukus):

\[x={{3y-5}\over{6y-1}}\] \[(6y-1)\times x=(6y-1)\times{{3y-5}\over{6y-1}}\] \[(6y-1)\times x=3y-5\]

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

6xy-x=3y-5 now what?

OpenStudy (unklerhaukus):

actually was it plus five or minus?

OpenStudy (anonymous):

+5

OpenStudy (anonymous):

what do i do from\[6xy-x=3y+5\]

OpenStudy (unklerhaukus):

minus five from both sides

OpenStudy (unklerhaukus):

minus 6xy from both sides

OpenStudy (anonymous):

\[-x-5=3y-6xy\]\[-x-5=y(-6x+3)\]\[{{-x-5}\over{-6x+3}}~~~=y\]\[y={{x+5}\over{6x-3}}\]

OpenStudy (anonymous):

you rock...i need to sleep...stay blessed bro....g'nite! :P

OpenStudy (unklerhaukus):

great work !